OpenStudy (anonymous):

Show that the number is rational by writing it as a quotient of two integers: -1 1/3

OpenStudy (anonymous):

thank you ^_^

OpenStudy (unklerhaukus):

ops i have made a mistake

OpenStudy (unklerhaukus):

\[-1\frac 1 3=-\frac 3 3 -\frac 1 3=\frac{-3-1}{3}=\frac{-4}3\]

OpenStudy (anonymous):

is that the right asnwer? hehe.

OpenStudy (unklerhaukus):

yes it is -4/3

OpenStudy (anonymous):

ok thank u~!